if the potential at infinity is -GM/r. then what is the potential at the surface of the earth. Please explain the approach

Grade:

2 Answers

vikas askiitian expert
13 years ago

potential at infinity is zero wrt earth...potential at surface is PE=gGM/r , then at inifinity it is zero...if potential at infinity is  -GM/r then it means -GM/r is added to its initial value so adding this at any point will give potential at that point...

so potential at the surface potential becomes GM/r-GM/r=0...

vikas askiitian expert
13 years ago

potential at surface of earth is -GM/r -GM/r =-2GM/r  not zero... coz actual potential at surface is -GM/r not  GM/r and by mistake i have written GM/r ....so dont get confuced...